OpenStudy (anonymous):

Suppose you got 8 mangoes and 3 apples for $18 and 3 mangoes and 5 apples for $14.50. How much does each mango and each apple cost? (I already know both equations but need help solving for M.)

OpenStudy (anonymous):

Please look at my question.

OpenStudy (anonymous):

Well, what are your equations?

OpenStudy (anonymous):

8m + 3a= 18 and 3m + 5a= 14.50

OpenStudy (anonymous):

now, via substiution or elimination some for M and A

OpenStudy (anonymous):

think of what number can the entire first equation be multiplied by and the same for the next to cancel (or eliminate) M

OpenStudy (anonymous):

I don't know which variable in which equation to sole for first.

OpenStudy (anonymous):

doesn't matter - either one is fine...

OpenStudy (anonymous):

multiply first equation by -3 and the 2nd one by 8 in order to cancel out the M

OpenStudy (anonymous):

I need help. I don't know what to do.

OpenStudy (anonymous):

8m + 3a= 18 and 3m + 5a= 14.50 Are your equation right? So what we do is solve the first one for a, okay? So first isolate a and get: 3a = -8m +18, then divide by 3, so you get: a = (8/3)m + 6. Now that you know a, put this in the second equation and get m. Once you have m, plug into 1 of two original equations to get the actual value of a.

OpenStudy (anonymous):

Alright I'll try.

OpenStudy (anonymous):

Does m= 1.5?

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

I got it. THANKSTO EVERYONE FOR YOUR HELP!

OpenStudy (anonymous):

Oh.